### v2

# AgentThink: A Unified Framework for Tool-Augmented Chain-of-Thought Reasoning in Vision-Language Models for Autonomous Driving

Abstract
Vision-Language Models (VLMs) show promise for autonomous driving, yet their struggle with hallucinations, inefficient reasoning, and limited real-world validation hinders accurate perception and robust step-by-step reasoning. To overcome this, we introduce AgentThink, a pioneering unified framework that, for the first time, integrates Chain-of-Thought (CoT) reasoning with dynamic, agent-style tool invocation for autonomous driving tasks. AgentThink’s core innovations include: (i) Structured Data Generation, by establishing an autonomous driving tool library to automatically construct structured, self-verified reasoning data explicitly incorporating tool usage for diverse driving scenarios; (ii) A Two-stage Training Pipeline, employing Supervised Fine-Tuning (SFT) with Group Relative Policy Optimization (GRPO) to equip VLMs with the capability for autonomous tool invocation; and (iii) Agent-style Tool-Usage Evaluation, introducing a novel multi-tool assessment protocol to rigorously evaluate the model’s tool invocation and utilization. Experiments on the DriveLMM-o1 benchmark demonstrate AgentThink significantly boosts overall reasoning scores by 53.91% and enhances answer accuracy by 33.54%, while markedly improving reasoning quality and consistency. Furthermore, ablation studies and robust zero-shot/few-shot generalization experiments across various benchmarks underscore its powerful capabilities. These findings highlight a promising trajectory for developing trustworthy and tool-aware autonomous driving models.
footnotetext: Equal contribution.footnotetext: Corresponding authors.footnotetext: This work was completed during the internships of Kangan Qian and Ziang Luo in Xiaomi Corporation.
1 Introduction
Recent advances in foundation models have opened new opportunities for autonomous driving, where pretrained Large Language Models (LLMs) [10, 2] and Vision-Language Models (VLMs) [35, 47, 28] are increasingly employed to enable high-level scene understanding, commonsense reasoning, and decision making. These models aim to transcend traditional perception pipelines—which rely on hand-crafted components such as object detection [20, 32], motion prediction [29, 40], and rule-based planning [5]—by providing richer semantic representations and broader generalization grounded in web-scale knowledge.
Many recent approaches recast autonomous driving tasks as visual question answering (VQA) problems, applying supervised fine-tuning (SFT) to foundation VLMs with task-specific prompts for object identification, risk prediction, or motion planning [33, 25, 44, 8, 36]. However, as shown in Fig.2 (a), these models typically treat reasoning as a static input-to-output mapping, neglecting the uncertainty, complexity, and verifiability essential to real-world decision making. Consequently, they often suffer from poor generalization, hallucinated outputs, and limited interpretability [42].
To improve robustness and transparency, recent works have explored incorporating chain-of-thought (CoT) reasoning into VLMs, as shown in Fig. 2 (b). Some approaches adopt rigid CoT templates [35, 14], promoting structured logic at the expense of flexibility. Others use open-ended reasoning formats [26, 15], but may overfit to token patterns and exhibit shallow or redundant reasoning. Moreover, most existing methods rely purely on imitation learning from curated trajectories, lacking the ability to detect knowledge uncertainty or invoke tools for intermediate verification [46].
Figure 1: The performance of proposed AgentThink framework on the DriveLMM-o1 benchmark.
Figure 2: Illustration of the motivation and key highlights of our proposed framework. (a) Base VLMs use static input-output mapping with no reasoning, leading to low accuracy and frequent hallucinations. (b) VLM + CoT introduces structured reasoning, improving interpretability, but still suffers from inconsistencies and lack of verification. (c) AgentThink (Ours) augments CoT with dynamic tool use, enhancing accuracy, reducing hallucinations, and improving reasoning consistency through external verification.
These challenges lead to a pivotal question: How can a VLM truly function as a decision-making agent—cognizant of its knowledge boundaries, proficient in verification, and capable of learning from tool-guided feedback? Inspiration comes from experienced human drivers who will consult aids like mirrors or GPS to refine their judgment when uncertain. Similarly, a capable autonomous agent must not only reason explicitly but also recognize its limitations and dynamically employ tools, such as object detectors or motion predictors, to steer its reasoning and decision-making.
Therefore, we present AgentThink, a unified framework for VLMs in autonomous driving that models reasoning not as a static output, but as an agent-style process—in which the model learns to utilize tools to generate Tool-Augmented reasoning chains, verify intermediate steps, and refine its conclusions. As illustrated in Fig.2 (c), rather than blindly mapping inputs to outputs, AgentThink dynamically decides when and how to use tools during inference to support or revise reasoning paths. To enable this behavior, we create a data-training-evaluation pipeline. First, we construct a structured dataset of Tool-Augmented reasoning traces. Then, we introduce a two-stage training pipeline: (i) SFT to warm up reasoning capabilities, and (ii) GRPO [31], a reinforcement learning (RL)-based strategy that refines reasoning depth and tool-use behavior through structured rewards. Finally, we propose a comprehensive evaluation protocol beyond answer correctness to assess tool selection, integration quality, and reasoning-tool alignment.
As shown in Fig. 1, experiments on the advanced DriveLMM-o1 benchmark [15] demonstrate that AgentThink achieves new state-of-the-art performance in both answer accuracy and reasoning score, surpassing existing models. The effectiveness of our approach in cultivating dynamic, tool-aware reasoning is further substantiated by comprehensive ablation studies and robust generalization capabilities across multiple benchmarks. These collective results strongly suggest that empowering vision-language agents with learned, dynamically invoked tool use is pivotal for creating more robust, interpretable, and generalizable systems for autonomous driving.
Generally, our contributions are as follows:
  • We propose AgentThink, the first framework to integrate dynamic, agent-style tool invocation into vision-language reasoning for autonomous driving tasks.
  • We develop a scalable data generation pipeline that produces structured, self-verified data with integrated tool usage and reasoning chains.
  • We introduce a two-stage training pipeline that combines SFT with GRPO, enabling models to learn when and how to invoke tools to enhance reasoning performance.
  • We design new evaluation metrics tailored to autonomous driving tool invocation, capturing tool selection, integration quality, and reasoning-tool alignment.
2 Related Works
2.1 Language Models in Autonomous Driving
Recent advancements in language modeling have opened up new opportunities for autonomous driving, particularly in enabling interpretable reasoning, commonsense understanding, and decision-making [6]. Early efforts integrated LLMs such as GPT series [27] by recasting driving tasks—e.g., scene description [44, 23], decision-making [9, 41], and risk prediction [3, 22]—into textual prompts, allowing for zero-shot or few-shot inference. While these approaches showcased the reasoning potential of LLMs, they often lacked step-by-step interpretability and struggled with generalization in out-of-distribution scenarios [39].
Recent works have augmented LLMs with prompting strategies, memory-based context construction, or vision inputs [12]. For instance, DriveVLM [35] introduces a CoT approach with modules for scene description, analysis, and hierarchical planning, while DriveLM [33] focuses on graph-structured visual question answering. EMMA [13] demonstrates how multimodal models can directly map raw camera inputs to driving outputs, including trajectories and perception objects. Despite these advancements, both LLM-centric and VLM-based methods often treat reasoning as a static input-output mapping, with limited ability to detect uncertainty, perform intermediate verification, or incorporate physical constraints [15]. Challenges such as hallucinations, over-reliance on rigid templates, and a lack of domain-specific reward feedback persist. To address these limitations, our work introduces a Tool-Augmented, RL-based reasoning framework that enables dynamic and verifiable decision-making for autonomous driving.
Figure 3: AgentThink Framework Architecture. (i) Structured and scalable data generation pipeline that constructs tool-augmented reasoning; (ii) Two-stage training pipeline that first performs SFT and then applies GRPO to improve reasoning and tool-use behavior; and (iii) Unified inference and evaluation protocol that dynamically invokes tools and assesses final answers based on reasoning completeness, consistency, and tool-use effectiveness.
2.2 Visual Question Answering in Autonomous Driving
Visual Question Answering (VQA) for autonomous driving has emerged as a benchmark paradigm to evaluate perception, prediction, and planning capabilities. Benchmarks such as BDD-X [18], DriveBench [42], DriveMLLM [11], Nuscenes-QA [30], and DriveLMM-o1 [15] provide structured QA tasks covering complex reasoning scenarios in urban and highway environments. For VQA tasks, recent approaches such as Reason2Drive [26], Alphadrive [17], OmniDrive [37], and DriveCoT [38] introduce COT reasoning to enhance model interpretability.
However, many adopt rigid reasoning templates or rely solely on imitation learning, making them prone to overfitting and hallucination. These methods often overlook dynamic reasoning processes and fail to verify intermediate steps using external tools. In contrast, our framework combines structured data generation with step-level rewards and tool-verification during inference. By employing RL via GRPO, we optimize the model’s reasoning trajectory to align with correctness, efficiency, and real-world applicability, setting a new direction for VQA in autonomous driving.